OpenStudy (anonymous):

PLEASE HELP WILL METAL Find the x value for point L such that JL and KL form a 3:1 ratio. Segment JK is shown. J is at negative 3, 4. K is at 4, negative 2. 5.25 2.33 2.25 4.33

OpenStudy (anonymous):

sorry.. i spelled medal wrong ha!

OpenStudy (campbell_st):

Ok... so this is how I solve this equation \[\frac{JL}{KL} = \frac{3}{1} \] cross multiply and \[JL = 3KL\] so (-3, 4) and 3 x (4, -2) so then \[x = \frac{1 \times -3 + 3 \times 4}{3 + 1}\] hope it helps

OpenStudy (anonymous):

i don't think i did this right. i got 2?

OpenStudy (campbell_st):

well I think you need to check the calculation. as if done correctly you get one of your answer choices.

OpenStudy (anonymous):

i don't think my calculator works for this type of equation..(i only have my phone calculator) :/

OpenStudy (anonymous):

is there another way i can find the answer?

OpenStudy (campbell_st):

well do it in bits... calculate the numerator.... then the denominator then to finish its numerator divided by the denominator... if that doesn't work use this web based calculator http://web2.0calc.com/

OpenStudy (campbell_st):

perhaps brackets around the numerator and brackets around the denominator with a divide symbol inbetween

OpenStudy (anonymous):

i got 4...

OpenStudy (anonymous):

is it 4.33?

OpenStudy (campbell_st):

wow... I'd give the phone away... so what is 1 x -3 = ? and what is 3 x 4 = ?

OpenStudy (anonymous):

yeah, seriously my phone is trash haha so 1x(-3)= 3 and 3x4= 12

OpenStudy (campbell_st):

wow... I think you need to check 1 x -3 =

OpenStudy (anonymous):

i meant -3.. my bad

OpenStudy (campbell_st):

so the numerator is now -3 + 12 = ?

OpenStudy (anonymous):

i got 9

OpenStudy (anonymous):

?

OpenStudy (campbell_st):

great so the denominator is 3 + 1 = 4 so what is 9/4 = ?

OpenStudy (anonymous):

ohh so its 2.25!

OpenStudy (campbell_st):

that's it...

OpenStudy (anonymous):

thanks a lot man :)

OpenStudy (campbell_st):

have fun